Cherry Blossom Products Inc. produces and sells yoga-training products: how-to DVDs and a basic equipment set (blocks, strap, and small pillows). Last year, Cherry Blossom Products sold 13,500 DVDs and 4,500 equipment sets. Information on the two products is as follows:

DVDs Equipment Sets
Price $8 $25
Variable cost per unit 4 15

Total fixed cost is $84,920.

Suppose that in the coming year, the company plans to produce an extra-thick yoga mat for sale to health clubs. The company estimates that 9,000 mats can be sold at a price of $15 and a variable cost per unit of $9. Total fixed cost must be increased by $28,980 (making total fixed cost $113,900). Assume that anticipated sales of the other products, as well as their prices and variable costs, remain the same.

Unless otherwise instructed, round all total dollar figures (e.g., sales, total contribution margin) to the nearest dollar, breakeven or target units to the nearest unit, and unit costs and unit contribution margins to the nearest cent. Round ratios to four significant digits.

1. What is the sales mix of DVDs, equipment sets, and yoga mats?
3:1:2

2. Compute the break-even quantity of each product.

Break-even DVDs units
Break-even equipment sets units
Break-even yoga mats units

3b. What is the overall contribution margin ratio? Use the contribution margin ratio to compute overall break-even sales revenue. (Note:Round the contribution margin ratio to two decimal places; round the break-even sales revenue to the nearest dollar.)

Overall contribution margin ratio 43.04 %
Overall break-even sales revenue $

4. Compute the margin of safety for the coming year in sales dollars.
$
Answer #1

Ans: Cherry Blossom Products Inc.

Basic Information

Particulars DVD Equipment sets Extra thick yoga mat
Sales (units) 13500 4500
Estimated sales next year (units) 9000
Sale Price($) 8 25 15
variable cost ($) 4 15 9
Contribution per unit ($) ( sale price - variable cost) 4 10 6
sales mix ratio 3 1 2

2. Break even sales in Quantity

Contribution per unit of sales mix= (3*4)+(1*10)+(2*6)\

= $ 34

Where one unit of sales mix contain 3 units of DVD, 1 unit of equipment set and 2 units of Extra thick yoga mat

Fixed cost = $113,900

Overall Break even sales per unit of sales mix = $113,900\ $ 34

= 3,350 units

So break even sales of

DVD = 3*3,350 = 10,050 units

Equipment set = 1*3,350= 3,350 units

Extra thick yoga mat= 2* 3,350 = 6,700 units

Point 3 b): Overall break even margin ratio= 43.04%

Overall break even sales = Fixed cost \ overall break even margin ratio

= $113,900\ 43.04%

= $264,638

Point 4): Margin of safety = Actual sales - break-even sales

Particulars DVD Equipment sets Extra thick yoga mat Total
Sales ( Units) 13,500 4,500 9,000
Sale price per unit ($) 8 25 15
Sales ($) 108,000 112,500 135,000 355,500
Break even sales 264,638
Margin of safety 90862

Margin of safety ratio = Margin of safety \ actual sales

= $ 90,862\ $355,500

= 0.2556 or 25.56%